# Phase transition to causal symmetry reveals operational autonomy in sociotechnical systems

### v1

@@ Abstract
A viability index, combining activity and structure, outperforms activity-based prediction ($\text{AUC}=0.88$ vs $0.81$), identifying ‘zombie’ systems where high churn masks structural decay.This extends recent work by Ait et al., who identified ‘zombie’ projects exhibiting activity without development based on non-coding contributions ait2022. Our metric identifies structural zombies: projects where coding activity persists but fails to preserve architectural invariants.
These results establish causal symmetrization as an empirically validated signature of self-organizing autonomy applicable across complex collaborative systems—a dynamical regime previously theorized in biological contexts but here demonstrated and measured in artificial ones.

### v2

# Causal symmetrization as an empirical signature of operational autonomy in complex systems

@@ Abstract
Theoretical biology has long proposed that autonomous systems sustain their identity through reciprocal constraints between structure and activity, a dynamical regime underlying concepts such as closure to efficient causation and autopoiesis. Despite their influence, these principles have resisted direct empirical assessment outside biological systems.
@@
A composite viability index combining activity and structural persistence outperforms activity-based prediction alone (AUC = 0.88 vs. 0.81), identifying “structural zombie” systems in which sustained activity masks architectural decay.
Together, these results show that causal symmetrization functions as a necessary statistical signature consistent with theoretical notions of operational closure, without implying biological life or mechanistic closure. They demonstrate that core principles of autonomy can be empirically probed in artificial collaborative systems, supporting substrate-independent dynamical signatures of self-organizing autonomy across complex adaptive systems.